Find the directional derivative of f(x,y,z)=z^3−x^(2)y at the point (-3, -4, 1) in the direction of the vector v=⟨4,−3,1⟩.
1
Expert's answer
2020-10-19T18:10:52-0400

"\nabla f=(-2xy,-x^2,3z^2)." Hence at (-3,-4,1) the value is (-24,-9,3). Unit vector along the direction is "\frac{(4,-3,1)}{\sqrt{}4^2+(-3)^2+1^2}=(\frac{4}{\sqrt{26}},\frac{-3}{\sqrt{26}},\frac{1}{\sqrt{26}})." Hence required derivative= "-24.\frac{4}{\sqrt{26}}-9.\frac{-3}{\sqrt{26}}+3.\frac{1}{\sqrt{26}}=\frac{-66}{\sqrt{26}}".